OpenStudy (anonymous):

What are the x-intercepts of the parabola x2 + 4x −12?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

(2, -16)

OpenStudy (anonymous):

so that means roots of this equation put y=0 x2+4x-12=0 (x+6)=0 and (x-2)=0 6 and 2

OpenStudy (anonymous):

x = -6 and x = 2

OpenStudy (anonymous):

a parabola which intersects the x axis has some solutions . and you know y is always 0 at x axis so to find the x intercepts we plug in y=0.

OpenStudy (precal):

 parabola with 2 complex solutions and dos not intersect the x-axis just a thought

OpenStudy (anonymous):

and we will say b2-4ac<0 :D